OpenStudy (anonymous):

-6= x/8 + 4 ... –16 1.8 16 –80

OpenStudy (anonymous):

\[\huge -6=\frac{x}{8}+4 ?\]?

OpenStudy (anonymous):

-6=x/8+4 so -4 from both sides -10=x/8 multiply both sides by 8 -80=x

OpenStudy (anonymous):

thank you!

OpenStudy (anonymous):

ur welcome